OpenStudy (anonymous):

what is the pH of a solution with a [H+]=3.74 x 10 ^-12 M?

OpenStudy (anonymous):

it would help to know what the formula for pH is. i believe it involves the common log

OpenStudy (anonymous):

yeah, pH= -log[H+] but i have trouble plugging it into my calculator.... which seems to be my biggest issue...

OpenStudy (anonymous):

Is that \(\log_{10}[H+]\)?

OpenStudy (anonymous):

Is it base 10?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Because \[ \log_{10}(10^{-12}) = -12 \]

OpenStudy (anonymous):

\[-\log(3.74\times 10^{-12})=-\log(3.74)+12\]

OpenStudy (anonymous):

If you suppose something like \(\sqrt{10}\approx 3.74\) then \(\log_{10}(3.74)\approx 1/2\)

OpenStudy (anonymous):

thx guys :)